# Dr.Occ: Depth- and Region-Guided 3D Occupancy from Surround-View Cameras for Autonomous Driving

### 01 Mar 2026 v1

@@ Abstract
footnotetext: Work done during internship at Horizon Robotics. #Project Leader.
@@ 3.1 Problem Formulation
We consider the task of vision-based 3D semantic occupancy prediction, where the system observes $T$ consecutive frames from $N_{c}$ calibrated surround-view cameras. For each camera $i\in\{1,\dots,N_{c}\}$, the intrinsic matrix is denoted as $\mathbf{K}_{i}\in\mathbb{R}^{3\times 3}$ and the extrinsic parameters as $\left[\mathbf{R}_{i}|\mathbf{t}_{i}\right]$, where $\mathbf{R}_{i}\in\mathbb{R}^{3\times 3}$ is the rotation matrix and $\mathbf{t}_{i}\in\mathbb{R}^{3}$ is the translation vector. At each time step $t\in[1,T]$, the multi-camera image set is represented as $\{\mathbf{I}_{1}^{(t)},\mathbf{I}_{2}^{(t)},\dots,\mathbf{I}_{N_{c}}^{(t)}\}$, where $\mathbf{I}_{i}^{(t)}\in\mathbb{R}^{H\times W\times 3}$ denotes an image of spatial resolution $H\times W$ from camera $i$. The goal is to predict a 3D voxel-wise semantic occupancy grid $\mathbf{O}\in\mathbb{R}^{X\times Y\times Z\times C}$ describing the environment around the ego-vehicle, where $(X,Y,Z)$ denote the spatial resolution and $C$ is the number of semantic classes. Following the prevailing paradigm [33, 8, 21], our pipeline first extracts 2D image features $\mathbf{F}_{2\mathrm{D}}^{(t)}\in\mathbb{R}^{H\times W\times D}$ from each frame $t\in[1,T]$ of the multi-camera input, then applies various transformers $\{\mathcal{T}\}$ to lift and aggregate features across all $T$ frames into a unified 3D feature volume $\mathbf{F}_{3\mathrm{D}}\in\mathbb{R}^{X\times Y\times Z\times D}$. Finally, a 3D decoder $\mathcal{D}$ maps $\mathbf{F}_{3\mathrm{D}}$ to the semantic occupancy prediction $\hat{\mathbf{O}}=\mathcal{D}(\mathbf{F}_{3\mathrm{D}})$.
@@ 3.5.2 Region-guided Expert Transformer
where $\mathcal{S}$ denotes the set of selected region indices and $K$ controls the number of activated experts. Each expert $E_{m}$ applies the same DCA module as in D2-VFormer, but restricted to its corresponding region via a binary spatial mask $\mathcal{M}_{m}$:
@@
where $w_{m}$ is the normalized weight for region $m$ computed from the importance scores, $E_{m}$ is the expert dedicated to region $\mathcal{R}_{m}$, and $\mathcal{M}_{m}$ is the binary spatial mask defining region $\mathcal{R}_{m}$. By dedicating experts to predefined regions and selectively activating them based on computed importance, R-EFormer concentrates model capacity on semantically important regions while efficiently handling spatial semantic variations.
@@ 3.5.3 Region-guided Recursive Expert Transformer
Let $\mathbf{F}^{(t)}$ denote the voxel features at iteration $t$ and $\mathcal{M}^{(t)}$ denote the spatial mask activated in iteration $t$. The mask $\mathcal{M}^{(t)}$ is generated by a router $\mathcal{R}^{(t)}$ that progressively focuses on the most salient regions. Specifically, the router first produces an importance weight map based on the input features and previous mask, from which the top-$k_{t}$ voxels are selected:
@@ 4.3 Ablation Study
The ablation results on Occ3D-nuScenes are summarized in Table 2. The baseline model achieves 70.36% IoU and 36.01% mIoU. Introducing the geometric enhancement module D2-VFormer yields clear gains of +0.93% IoU and +5.44% mIoU, verifying its effectiveness in geometric completion and semantic improvement. Combining D2-VFormer with the region-guided semantic decoder R-EFormer further raises performance to 73.45% IoU and 43.03% mIoU, highlighting the complementarity between geometric and semantic enhancement. Replacing R-EFormer with its recursive variant R2-EFormer slightly reduces IoU but achieves the highest mIoU of 43.43%, owing to its stronger focus on rare and hard-to-recognize categories through adaptive recursive refinement.

### 04 Mar 2026 v2

@@ Abstract
footnotetext: Work done during internship at Horizon Robotics. #Project Lead.
@@ 3.1 Problem Formulation
We consider the task of vision-based 3D semantic occupancy prediction, where the system observes $T$ consecutive frames from $N_{c}$ calibrated surround-view cameras. For each camera $i\in\{1,\dots,N_{c}\}$, the intrinsic matrix is denoted as $\mathbf{K}_{i}\in\mathbb{R}^{3\times 3}$ and the extrinsic parameters as $\left[\mathbf{R}_{i}|\mathbf{t}_{i}\right]$, where $\mathbf{R}_{i}\in\mathbb{R}^{3\times 3}$ is the rotation matrix and $\mathbf{t}_{i}\in\mathbb{R}^{3}$ is the translation vector. At each time step $t\in[1,T]$, the multi-camera image set is represented as $\{\mathbf{I}_{1}^{(t)},\mathbf{I}_{2}^{(t)},\dots,\mathbf{I}_{N_{c}}^{(t)}\}$, where $\mathbf{I}_{i}^{(t)}\in\mathbb{R}^{H\times W\times 3}$ denotes an image of spatial resolution $H\times W$ from camera ii. The goal is to predict a 3D voxel-wise semantic occupancy grid $\mathbf{O}\in\mathbb{R}^{X\times Y\times Z\times C}$ describing the environment around the ego-vehicle, where $(X,Y,Z)$ denote the spatial resolution and $C$ is the number of semantic classes. Following the prevailing paradigm [33, 8, 21], our pipeline first extracts 2D image features $\mathbf{F}_{2\mathrm{D}}^{(t)}\in\mathbb{R}^{H\times W\times D}$ from each frame $t\in[1,T]$ of the multi-camera input, then applies various transformers $\{\mathcal{T}\}$ to lift and aggregate features across all $T$ frames into a unified 3D feature volume $\mathbf{F}_{3\mathrm{D}}\in\mathbb{R}^{X\times Y\times Z\times D}$. Finally, a 3D decoder $\mathcal{D}$ maps $\mathbf{F}_{3\mathrm{D}}$ to the semantic occupancy prediction $\hat{\mathbf{O}}=\mathcal{D}(\mathbf{F}_{3\mathrm{D}})$.
@@ 3.5.2 Region-guided Expert Transformer
where $\mathcal{S}$ denotes the set of selected region indices and $K$ controls the number of activated experts. Each expert $E_{m}$ applies the same DCA module as in D2-VFormer, but restricted to its corresponding region via a binary mask $\mathcal{M}_{m}$:
@@
where $w_{m}$ is the normalized weight for region $m$ computed from the importance scores, $E_{m}$ is the expert dedicated to region $\mathcal{R}_{m}$, and $\mathcal{M}_{m}$ is the binary spatial mask defining region $\mathcal{R}_{m}$. By dedicating experts to predefined regions and selectively activating them based on importance, R-EFormer concentrates model capacity on semantically important regions while handling spatial semantic variations.
@@ 3.5.3 Region-guided Recursive Expert Transformer
Let $\mathbf{F}^{(t)}$ denote the voxel features at iteration $t$ and $\mathcal{M}^{(t)}$ denote the spatial mask in iteration $t$. The mask $\mathcal{M}^{(t)}$ is generated by a router $\mathcal{R}^{(t)}$ that progressively focuses on the most salient regions. Specifically, the router first produces an importance weight map based on the input features and previous mask, from which the top-$k_{t}$ voxels are selected:
@@ 4.3 Ablation Study
The ablation results on Occ3D-nuScenes are summarized in Table 2. The baseline model achieves 70.36% IoU and 36.01% mIoU. Introducing the geometric enhancement module D2-VFormer yields clear gains of +0.93% IoU and +5.44% mIoU, verifying its effectiveness in geometric completion and semantic improvement. Combining D2-VFormer with the semantic decoder R-EFormer further raises performance to 73.45% IoU and 43.03% mIoU, highlighting the complementarity between geometric and semantic enhancement. Replacing R-EFormer with its recursive variant R2-EFormer slightly reduces IoU but achieves the highest mIoU (43.43%), owing to its stronger focus on rare and hard-to-recognize categories through adaptive recursive refinement.
@@ 5 Conclusion
Acknowledgement
The work was supported by NSFC (62301370), GuangDong Basic and Applied Basic Research Foundation (2025A1515010440), and Shenzhen Science and Technology Program (JCYJ20240813111202004).